I need someone to please review my assignment solutions...I would greatly appreciate it....I was supposed to solve for (x-5)^2=6 MY ANSWER:5+sqrt of 6....Next problem is 5(x-2)^2=3 MY ANSWER:10+sqrt of 15 over 5 Am I correct with my solutions??? If not could someone please help in explainimh where I went wrong...Thanks ;0)
Answer by stanbon(75887)

(x-5)^2=6 MY ANSWER:5+sqrt of 6....Next problem is 5(x-2)^2=3 MY ANSWER:10+sqrt of 15 over 5
-----------
You have one of the two solutions for each problem. Quadratics have two solutions>
--------
(x-5)^2=6
Take the square root to get:
x-5 = sqrt6 or x-5 = -sqrt6
x= 5+sqrt6 or x=5-sqrt6
----------------------------
5(x-2)^2=3
(x-2)^2 = 3/5
Take the square root to get:
x-2 = sqrt(3/5) = - + [sqrt15]/5
x= 2+[sqrt(15)]/5 or x = 2 -[sqrt(15)]/5
==============
Cheers,
Stan H.